# Denoising Vision Transformers

### v1

@@ 6 Discussion and Future Works
Our work has introduced DVT, a robust method leveraging neural fields to eliminate feature artifacts from ViTs. We pinpoint positional embeddings as the primary source of these artifacts, despite their importance in various vision tasks. Utilizing a neural-field optimization process, DVT efficiently extracts clean features from the noise-riddled feature maps of existing ViTs. Building upon this, we propose a scalable feature denoiser, eliminating the need for individual image optimizations. When learned from a few denoised samples, our denoiser generalizes well to unseen data, and improves pre-trained ViTs by large margins in dense vision tasks. Furthermore, our research suggests several avenues for future exploration: Understanding the role of positional embeddings in ViTs could inform the design of next-generation deep learning architectures. Redefining positional embeddings within ViTs and Transformers is also an imperative problem. Finally, devising a method to denoise pre-trained ViT features without additional training presents a fascinating challenge.
@@ Acknowledgements
We are grateful to many friends, including Congyue Deng, Jiageng Mao, Junjie Ye Justin Lovelace, Varsha Kishore, and Christian Belardi, for their fruitful discussions on this work and follow-ups. We acknowledge an unrestricted gift from Google in support of this project.
Supplementary Material
In the appendix, we include comprehensive implementation details (§A) as well as discussions on the understanding of ViTs (§B), focusing specifically on the nuances of position embeddings. Following this, we discuss the limitations of this work and propose avenues for future exploration (§C).
@@ A.1 Denosing with Neural Fields
Recall that we decompose the output feature map from a pre-trained ViT into three components: $\mathbf{y}\approx\mathcal{F}(\mathcal{A})+\mathcal{G}+\mathbf{h}(\mathbf{y})$, where $\mathcal{F}$ is a s semantic field, $\mathcal{G}$ is an artifact field, and $\mathbf{h}$ is a residual predictor. We describe their implementation details below.
@@
To facilitate efficient learning, we use InstantNGP [20], a type of compact and fast coordinate network, parameterized by learnable multi-level hash grids $\mathcal{H}$ and a lightweight MLP $\phi(\cdot)$, to learn $\mathcal{F}$. It takes as input a normalized 2D coordinate $(i,j)$, within the range of [0, 1], and outputs its corresponding feature vector, i.e., $\mathcal{F}(i,j)=\phi\left(\mathcal{H}(i,j)\right)$. We refer readers to [20] for a more detailed understanding of the learnable hash grids. In our implementation, we use a hash encoding resolution that spans from $2^{4}$ to $2^{10}$ with 16 levels. Each hash entry has a channel size of 8. The maximum number of hash entries of each resolution is $2^{20}$. For the lightweight MLP, we use a two-layer Linear-ReLu-Linear structure. The hidden dimension of this MLP is half the size of the output feature dimension, and the output feature dimension itself corresponds to the feature dimension of the ViT being studied (e.g. 768 for a ViT-B and 1024 for a ViT-L).
Figure S6: Feature map visualizations: positional embeddings (PE) and a cat image in different ViTs. We visualize the feature maps across different layers (1 to 12) of various pre-trained ViT-base models, displayed sequentially from left to right. For each panel, the top row shows the feature maps generated by inputting zero-tensors, highlighting the influence of PE alone. The middle row showcases the feature norm of the PE feature map. The bottom row presents the feature map for a sample cat image, allowing for a comparison that reveals visual correlations between the artifacts in general image feature maps and the PE feature map.
@@
For all experiments, we use a 2D learnable feature map of size $C\times K\times K$ to learn the input-independent noise, where $C$ corresponds to the feature dimension of the studied ViT, and $K$ is the spatial size. We compute $K$ by $(H-P)/S+1$, where $H$ is the height&width of input images (which we resize to be square), $P$ is the patch size, and $S$ is the stride size used in the model. To accommodate ViTs with different patch sizes, we set $H$ to 518 for those trained with a patch size of 14, and 512 for ViTs with a patch size of 16, resulting in $K$ values of 37 and 32, respectively. Note that this feature map, $\mathcal{G}$, can be interpolated to fit any arbitrary image size. We specifically choose these $K$ values to minimize the need for interpolation during training, thus enhancing training efficiency.
@@
In our implementation, we extract $N=768$ crops from each image, applying random augmentations, which include random flipping with a probability of 0.5, and random resizing and cropping, where the size of the crop is scaled between 0.1 to 0.5 of the original image size and the aspect ratio is maintained between 3/4 and 4/3.
The coefficients in our loss function (Equation 6) are set as $\alpha=0.1$ and $\beta=0.02$. We use Adam optimizer, with a learning rate of 0.01 and a LinearLR decay strategy. Our models are trained for 20,000 iterations. Each iteration will process 2048 randomly sampled pixels from the pre-extracted feature maps. Note that due to the efficient implementation of $\mathcal{F}$ and the pre-extraction of patch features, our denoising typically takes about 100-160 seconds to finish (including the feature extraction time). This rapid optimization process allows us to easily amortize the denoising cost with parallel computes, thereby ensuring the practicality and applicability of our method in various scenarios.
We use the same hyperparameters for all experiments without any specific tuning. See Figures S9, S10, S12, S11, S15, S14 and S13 for visualizations of some examples of our per-image denoising outputs.
@@ A.2 Generalizable Denoiser
Optimization
To train the denoiser, we optimize the loss function defined in Equation 11. Note that our approach does not necessitate re-training ViTs; instead, it only optimizes the newly initialized parameters. The denoiser is trained over 10 epochs with a batch size of 64, utilizing the AdamW optimizer with a learning rate of 2e-4 and a cosine learning rate scheduler. The denoiser training typically takes about 2 hours on 8 GPUs.
@@ A.3 ViT Models
We provide the timm model identifiers of the ViTs studied in this paper in Table S5. For experiments with large input image sizes (e.g. using the 512-sized images as input to a model trained with 224-image-resolution), we always resize the position embeddings using bicubic interpolation to accommodate the increased size.
Table S5: timm model indentifiers.
@@ A.4 Correlation
In the main text, we mention the correlation between artifacts and their positions in images without detailed context, which we now provide. Our focus is on quantifying the correlation between different features and their positions within an image. To analyze this correlation, we employ the maximal information coefficient (MIC), a metric originally used for measuring the strength of linear or non-linear associations between two scalar variables. To adapt MIC for our purpose, we compute the association between high-dimensional features $\mathbf{f}$ and their positions. We calculate this by taking the maximal MIC across all channels of $\mathbf{f}$ and averaging the MICs of the $x$ and $y$ coordinates:
@@
Our analysis includes a comparison of MIC values for the decomposed noise map, the original noisy ViT features, and the denoised, artifact-free features. The results, present in Table 1 of the main paper, reveal that the decomposed noise map exhibits the highest correlation with image positions. The noisy features, entangled with noise artifacts originating from the position embeddings, display the second-highest positional correlation. In contrast, the noise-free features denoised by our method show the lowest correlation with positions, demonstrating the effectiveness of our decomposition approach in removing such artifacts.
@@ A.5 Feature Qualitative Results
We additionally visualize the features for algorithms with weak artifacts in Figure S8. We empirically observe that ViTs trained using both MAE and DINO exhibit very few visible artifacts in their feature (center column). Figures S15 and S14 show additional visualizations of the decomposed noise map and the learned residual terms of MAE and DINO, respectively. We note that decomposed noise maps from these two models typically manifest low-frequency patterns and the residual terms do not yield pronounced patterns.
@@
Additional visualizations of the feature maps at all layers of ViT models are shown in Figure S6. Observe that the artifact is present in almost all layers of the models. See Figures S9, S10, S12, S11, S15, S14 and S13 for more visualizations.
Appendix B Further Discussion into ViT Understanding
High-norm vs. Low-norm patterns.
The concurrent research [8] identifies artifacts in ViTs by examining the feature norm. However, our findings, as illustrated in Figure 5 “Original Feature L2 Norm” columns (e.g., there are many empty dark patches in the DEiT-III visualization (c)), reveals the existence of “low-norm” patterns that also behave as artifacts, especially in models like CLIP [26] and DeiT-III [32]. In addition, the research [8] concludes that the “high-norm” patterns are particularly pronounced in large and sufficiently trained ViTs, a trend that is not observed in small models, but our analysis showcases the presence of artifacts in almost all ViTs. These discoveries suggest that solely assessing artifacts based on feature norms does not provide a comprehensive understanding. Consequently, this calls for more in-depth research to fully understand the nature of artifacts in ViTs and how they impact model performance and behavior. Such insights are crucial for developing more robust, next-generation ViTs, particularly in the context of handling and interpreting features in ViTs.
@@ Appendix B Further Discussion into ViT Understanding
A key component of our hypothesis as to why artifacts exist in ViT features is the use of positional embeddings. Currently, all ViTs leverage either fixed [15] or learned [3, 22, 32] positional embeddings that are added to the input tokens of the Transformer model. Alternatively, Rotary Positional Embeddings [30], which were originally proposed in the language domain for better sequence length scaling, does not directly add anything to the input tokens. Instead, this method encodes the absolute position with a rotation matrix and crucially incorporates the explicit relative position dependency in the computation of the attention values. While EVA02 [13] does leverage this kind of positional embedding, the training process involves distilling from the already-noisy features from CLIP. Indeed, the noisy artifacts of the EVA02 model bear semblance to those from CLIP models, especially in the later layers (Figure S6). Thus, while the positional embedding selection is promising, more research should be done towards ViTs that leverage these Rotary PE for artifact reduction. Similarly, the positional embedding used in the T5 language model [27] does not add a positional embedding directly to the input; instead, it learns a bias that is added to the key-query dot-product in the self-attention step and does not include explicit position information into the self-attention value vectors. ALiBi [23], used in many large language models (LLM), also does not do so, and instead adds a static bias to the query-key dot product. These methods eliminate the input-independent portion of the final output feature while retaining the benefits of the position embedding. For future work, we suggest further exploration into adapting other such positional embedding paradigms specifically for the image domain.
@@
Figure S9: Visualization of DINOv2 [22] per-image denoising. We visualize (a) the input image, (b) the similarity between the central red patch and other patches, (c) the original noisy feature map, (d) the denoised feature map, and (e) the similarity post-denoising. Additionally, we show the (f) decomposed noise map $\mathcal{G}$ and (g) its L2 norm as well as (h) the L2 norm of the predicted residual term $\mathbf{h}$.
@@
Figure S10: Visualization of CLIP [26] per-image denoising. We visualize (a) the input image, (b) the similarity between the central red patch and other patches, (c) the original noisy feature map, (d) the denoised feature map, and (e) the similarity post-denoising. Additionally, we show the (f) decomposed noise map $\mathcal{G}$ and (g) its L2 norm as well as (h) the L2 norm of the predicted residual term $\mathbf{h}$.
@@
Figure S11: Visualization of EVA02 [13] per-image denoising. We visualize (a) the input image, (b) the similarity between the central red patch and other patches, (c) the original noisy feature map, (d) the denoised feature map, and (e) the similarity post-denoising. Additionally, we show the (f) decomposed noise map $\mathcal{G}$ and (g) its L2 norm as well as (h) the L2 norm of the predicted residual term $\mathbf{h}$.
@@
Figure S12: Visualization of DeiT-III [32] per-image denoising. We visualize (a) the input image, (b) the similarity between the central red patch and other patches, (c) the original noisy feature map, (d) the denoised feature map, and (e) the similarity post-denoising. Additionally, we show the (f) decomposed noise map $\mathcal{G}$ and (g) its L2 norm as well as (h) the L2 norm of the predicted residual term $\mathbf{h}$.
@@
Figure S13: Visualization of DINOv2 with Registers [8] per-image denoising. We visualize (a) the input image, (b) the similarity between the central red patch and other patches, (c) the original noisy feature map, (d) the denoised feature map, and (e) the similarity post-denoising. Additionally, we show the (f) decomposed noise map $\mathcal{G}$ and (g) its L2 norm as well as (h) the L2 norm of the predicted residual term $\mathbf{h}$.
@@
Figure S14: Visualization of DINO [4] per-image denoising. We visualize (a) the input image, (b) the similarity between the central red patch and other patches, (c) the original noisy feature map, (d) the denoised feature map, and (e) the similarity post-denoising. Additionally, we show the (f) decomposed noise map $\mathcal{G}$ and (g) its L2 norm as well as (h) the L2 norm of the predicted residual term $\mathbf{h}$.
@@
Figure S15: Visualization of MAE [15] per-image denoising. We visualize (a) the input image, (b) the similarity between the central red patch and other patches, (c) the original noisy feature map, (d) the denoised feature map, and (e) the similarity post-denoising. Additionally, we show the (f) decomposed noise map $\mathcal{G}$ and (g) its L2 norm as well as (h) the L2 norm of the predicted residual term $\mathbf{h}$.
Appendix C Discussion on Limitations

### v2

@@ 6 Discussion and Future Works
Denoising Vision Transformers (DVT) introduces a robust method leveraging neural fields to eliminate feature artifacts from ViTs. This work additionally pinpoint positional embeddings as the primary source of these artifacts, despite their importance in various vision tasks. Using a neural field optimization process, DVT efficiently extracts clean features from the noise-riddled feature maps of existing ViTs. And using a scalable feature denoiser model, DVT eliminates the need for individual image optimizations. When learned from individually denoised samples, our denoiser generalizes well to unseen data and improves pre-trained ViTs by large margins in dense vision tasks. More broadly, our research suggests several avenues for future exploration: (1) understanding the role of positional embeddings in ViT could inform the design of next-generation deep learning architectures, and (2) redefining positional embeddings within ViTs and transformers is also an imperative problem. Lastly, combining the insights from our work and those of [7] could lead to a more complete picture of how these artifacts emerge. We hope that the results presented in this work contribute to a deeper understanding of artifacts in vision transformers and beyond.
@@ Acknowledgements
We are grateful to many friends, including Jiageng Mao, Junjie Ye, Justin Lovelace, Varsha Kishore, and Christian Belardi, for their fruitful discussions on this work and follow-ups. Katie Luo is supported by an Nvidia Graduate Fellowship. Leonidas Guibas acknowledges the support from a Vannevar Bush Faculty Fellowship. We also acknowledge an unrestricted gift from Google in support of this project.
Supplementary Material:
Denoising Vision Transformers
In the appendix, we provide detailed implementation details in §A, elaborate on evaluation protocols and additional results in §B, and discuss the understanding of position embeddings in ViT in §C. Lastly, we discuss the limitations of this work and propose a few avenues for future work in §D.
@@ A.1 Denosing with Neural Fields
Recall that we decompose the output feature map from a pre-trained ViT into three components: $\mathbf{y}\approx\mathcal{F}(\mathcal{A})+\mathcal{G}+\mathbf{h}(\mathbf{y})$, where $\mathcal{F}$ is a feature semantic field, $\mathcal{G}$ is an artifact field, and $\mathbf{h}$ is a residual predictor. We describe their implementation details below.
@@
To facilitate efficient learning, we use InstantNGP [23], a type of compact and fast coordinate network, parameterized by learnable multi-level hash grids $\mathcal{H}$ and a lightweight MLP $\phi(\cdot)$, to learn $\mathcal{F}$. It takes as input a normalized 2D coordinate $(i,j)$, within the range of [0, 1], and outputs its corresponding feature vector, i.e., $\mathcal{F}(i,j)=\phi\left(\mathcal{H}(i,j)\right)$. We refer readers to [23] for a more detailed understanding of the learnable hash grids. In our implementation, we use a hash encoding resolution that spans from $2^{4}$ to $2^{10}$ with 16 levels. Each hash entry has a channel size of 8. The maximum number of hash entries of each resolution is $2^{20}$. For the lightweight MLP, we use a two-layer Linear-ReLu-Linear structure. The hidden dimension of this MLP is half the size of the output feature dimension, which corresponds to the feature dimension of the ViT being studied (e.g., 768 for a ViT-base and 1024 for a ViT-large).
Figure S1: Feature map visualizations: positional embeddings (PE) and a cat image in different ViTs. We visualize the feature maps across different layers (1 to 12) of various pre-trained ViT-base models, displayed sequentially from left to right. For each panel, the top row shows the feature maps generated by inputting a zero tensor, highlighting the influence of PE alone. The middle row presents the feature norm of the PE feature map. The bottom row presents the feature map for a sample cat image, allowing for a comparison that reveals visual correlations between the artifacts in general image feature maps and the PE feature map.
@@
For all experiments, we use a 2D learnable feature map of size $C\times K\times K$ to learn the input-independent noise, where $C$ corresponds to the feature dimension of the studied ViT, and $K$ is the spatial size. We compute $K$ by $(H-P)/S+1$, where $H$ is the height&width of input images (which we resize to be square), $P$ is the patch size, and $S$ is the stride size used in the model. To accommodate ViTs with different patch sizes, we set $H$ to 518 for those trained with a patch size of 14, and 512 for ViTs with a patch size of 16, resulting in $K$ values of 37 and 32, respectively. Note that this feature map, $\mathcal{G}$, can be bilinearly interpolated to fit any arbitrary image resolution. We specifically choose these $K$ values to minimize the need for run-time interpolation during training, thus improving denoising efficiency.
@@
In our implementation, we extract $N=768$ views (crops) from each image, applying random augmentations, which include random flipping with a probability of 0.5, and random resizing and cropping, where the size of the crop is scaled between 0.1 to 0.5 of the original image size and the aspect ratio is maintained between 3/4 and 4/3. For understanding how the number of views used for training affects the DVT’s performance, please refer to Fig. 7 in the main text (our default setting is $N=768$).
The coefficients in our loss function ((Eq. 6) of the main text) are set as $\alpha=0.1$ and $\beta=0.02$. We use Adam optimizer, with a learning rate of 0.01 and a LinearLR decay strategy. Our models are trained for 20,000 iterations. Each iteration will process 2048 randomly sampled pixels from the pre-extracted feature maps. Note that due to the efficient implementation of $\mathcal{F}$ and the pre-extraction of patch features, our denoising typically takes about 100-160 seconds to finish (including the feature extraction time). This rapid optimization process allows us to easily amortize the denoising cost with parallel computes, thereby ensuring the practicality and applicability of our method in various scenarios.
We use the same hyperparameters for all experiments without any specific tuning. See Figs. S4, S5, S7, S6, S10, S9 and S8 for visualizations of some examples of our per-image denoising output.
@@ A.2 Generalizable Denoiser
Optimization.
To train the denoiser, we optimize the loss function defined in Eq. 11 of the main text. Note that our approach does not necessitate re-training ViTs; instead, it only optimizes the smaller denoisier network, which constitutes only 8% of the original model’s size. The denoiser is trained for 10 epochs with a batch size of 64, using the AdamW optimizer with a learning rate of $2e-$4 and a cosine learning rate scheduler. The denoiser training typically takes about 2 hours on 8 GPUs.
@@ A.3 ViT Models
We provide the timm model identifiers of the ViTs studied in this paper in Tab. S1. For experiments with large input image sizes (e.g. using the 512-sized images as input to a model trained with 224-image resolution), we always resize the position embeddings using bicubic interpolation to accommodate the increased size.
Table S1: timm model identifiers
@@ A.4 Correlation
In the main text, we mention the correlation between artifacts and their positions in images without a detailed context, which we now provide. Our focus is on quantifying the correlation between different features and their positions within an image. To analyze this correlation, we employ the maximal information coefficient (MIC), a metric originally used for measuring the strength of linear or nonlinear associations between two scalar variables. To adapt MIC for our purpose, we compute the association between high-dimensional features $\mathbf{f}$ and their positions. We calculate this by taking the maximal MIC across all channels of $\mathbf{f}$ and averaging the MICs of the coordinates $x$ and $y$.
@@
Our analysis includes a comparison of the MIC values for the decomposed noise map, the original noisy ViT features, and the denoised, artifact-free features. The results, presented in Tab. 1 of the main paper, reveal that the decomposed noise map exhibits the highest correlation with image positions. The noisy features, which are entangled with noise artifacts originating from the position embeddings, display the second highest positional correlation. In contrast, the noise-free features denoised by our method show the lowest correlation with positions, demonstrating the effectiveness of our decomposition approach in removing such artifacts.
@@ A.5 Feature Qualitative Results
We additionally visualize the features for algorithms with weak artifacts in Fig. S3. We empirically observe that ViTs trained using both MAE and DINO exhibit very few visible artifacts in their feature (center column). Figs. S9 and S10 shows additional visualizations of the decomposed noise map and the learned residual terms of DINO and MAE, respectively. We note that decomposed noise maps from these two models typically manifest low-frequency patterns and the residual terms do not yield pronounced patterns.
@@
Additional visualizations of the feature maps at all layers of ViT models are shown in Fig. S1. Observe that the artifact is present in almost all layers of the models. See Figs. S4, S5, S7, S6, S10, S9 and S8 for more visualizations.
Appendix B Evaluation Protocols
We introduce our evaluation protocols here, mostly following [25]. We have released our code, checkpoints, and logs for reproducibility at https://jiawei-yang.github.io/DenoisingViT/.
Semantic Segmentation.
We use a linear evaluation setting. In detail, we extract the final feature maps from the frozen backbone and pass them through the denoisers if there are any. Following this, feature maps are resized back to their original resolution. Then, a single learnable linear layer is trained to predict the semantic segmentation from these resized feature maps. The training and testing image resolutions are $518\times 518$, following [25]. We train this linear head for $40,000$ iterations for both VOC and ADE20k datasets. We report the mean intersection over union (mIoU) metric for all experiments.
Depth estimation.
We extract the final feature maps from the frozen backbone and pass them through the denoisers, if applicable. Then, we follow [25] to append the cls token to every patch token to enrich feature representations. We bilinearly upsample these features by a factor of 4 and train a linear layer using classification loss to divide the prediction into 256 uniformly distributed bins. Unlike [25], we slightly decrease the learning rate from 1e-4 to 5e-3, as we find that this modification improves most of the methods, including baselines, in our early experiments. We report our results on the commonly used metrics: AbsRel (absolute relative error |$\absolutevalue{d^{*}-d}/d$) and $\delta_{1}$ (percentage of pixels where $\max(d^{*}/d,~{}d/d^{*})<1.25$).
Object detection.
To evaluate the object detection task, we utilize the ViTDet detector [21] to infer object bounding boxes based on feature maps extracted either from original ViTs or denoisers. The detection framework is FasterRCNN [31]. The input image resolution for training and testing is $518\times 518$, the same as the semantic segmentation task. We train all models for 24k iterations, where we decay the learning rate at steps 20k and 22k.
Our initial attempts at directly learning an object detection head from the denoised features did not achieve superior performance. This led us to speculate that the omission of relative positional information, which was largely removed during denoising, might be important for accurately predicting the relative box coordinates of objects within the full image context. This requirement is almost unique to the bounding box prediction task. To counteract this, we re-add fixed sinusoidal positional embeddings into the feature maps produced by the denoisers. This adjustment, adding no additional learnable parameters, is found to enhance the detection performance. We believe that the disentanglement between positional features and semantic features would be an interesting direction to study. We also apply this method to the baseline models, and the results are shown in Tab. S2. We see that adding this step to the baselines does not yield consistent performance gains. Consequently, we apply this step only to our denoisers.
Table S2: Object detection with frozen features. We report the mAP metric on the VOC object detection benchmark. “fixed PE”: fixed sinusoidal positional embeddings.
Method DINOv2 DINOv2-reg DeiT-III CLIP DINO EVA02 Avg
baseline 81.4 80.9 80.9 75.8 76.4 79.4 79.2
+fixed PE 81.5 (+0.1) 81.2 (+0.3) 80.9 75.7 (-0.1) 75.8 (-0.6) 78.7 (-0.7) 79.0 (-0.2)
  +DVT 81.9 (+0.5) 81.4 (+0.5) 81.7 (+0.9) 77.0 (+1.2) 77.1 (+0.7) 80.2 (+0.8) 79.9 (+0.7)
Object discovery.
We use LOST [33] to evaluate the object discovery performance. LOST leverages the activation features of a pre-trained ViT for automated object discovery. Specifically, it uses the components of the last attention layer for computing the similarities between the different patches to discover and identify the object connected components. To use LOST, one has to manually sweep between query, key, value, or other intermediate model outputs as the indictor of objects’ prominence. Through our qualitative analysis, we find that the feature norm is a good candidate to indict object prominence (See Fig. 6). We report our results on the CorLoc metric (percentage of predicted box with an IoU greater than 0.5 with one of the labeled object bounding boxes) as in [33, 7].
Classification.
Although the global-level classification task is beyond the scope of our approach, our DVT demonstrates improved performance over its baselines through the use of an attentive probe protocol. Following the methodology described in AIM [10], we conduct an “attentive probe” on both the original and denoised patch tokens, omitting the CLS token, which our approach does not process during training. This probe employs attention mechanisms to maximize the extraction of information from each patch token. The backbones and the denoisers are frozen during our evaluation, and we train the attentive layer for 10 epochs. The results, presented in Tab. S3, suggest that DVT can potentially improve over its baselines, even though the denoising objective is orthogonal to classification. We believe integrating the CLS token into the denoising process represents a promising avenue for future research to enhance classification performance.
Additionally, we underscore the versatility of the denoiser in our DVT as a plug-in-and-play module, which can be optionally activated or deactivated to support various functionalities without compromising any properties of the original models. In essence, by leveraging the original class tokens before the denoiser, one can always recover the original models’ classification performance.
Table S3: ImageNet Classification Accuracy using Attentive Probing.
Method DINOv2 DINOv2-reg DeiT-III CLIP DINO EVA02
baseline 81.2% 81.6% 81.6% 83.4% 77.0% 80.3%
  +DVT 81.8% (+0.6) 81.8% (+0.2) 82.1% (+0.5) 83.5% (+0.1) 77.0% 80.5% (+0.2)
Appendix C Further Discussion into ViT Understanding
@@
A key component of our hypothesis as to why artifacts exist in ViT features is the use of positional embeddings. Currently, all ViTs leverage either fixed [16] or learned [4, 25, 36] positional embeddings that are added to the input tokens of the Transformer model. Alternatively, Rotary Positional Embeddings [34], which were originally proposed in the language domain for better sequence length scaling, does not directly add anything to the input tokens. Instead, this method encodes the absolute position with a rotation matrix and crucially incorporates the explicit relative position dependency in the computation of the attention values. Although EVA02 [13] does leverage this kind of positional embedding, the training process involves distilling from the already-noisy features of CLIP. Indeed, the noisy artifacts of the EVA02 model resemble those of CLIP models, especially in the later layers (Fig. S1). Thus, while the positional embedding selection is promising, more research should be done towards ViTs that leverage these Rotary PE for artifact reduction. Similarly, the positional embedding used in the T5 language model [30] does not add a positional embedding directly to the input; instead, it learns a bias that is added to the key-query dot product in the self-attention step and does not include explicit position information in the self-attention value vectors. ALiBi [26], used in many large language models (LLM), also does not do so, and instead adds a static bias to the query-key dot product. These methods eliminate the input-independent portion of the final output feature while retaining the benefits of the position embedding. For future work, we suggest further exploration into adapting other such positional embedding paradigms specifically for the image domain.
@@
Figure S4: Visualization of DINOv2 [25] per-image denoising. We visualize all components of the per-image denoising stage. From left to right: In the first 5 columns we visualize the input image, the original noisy feature map from the model, the K-Means clusters on the original features, the L2 norm on the original features, and the similarity between the central red patch and other patches. In the next 4 columns we visualize the the denoised feature map using DVT, the denoised features’ K-means clusters, the denoised features’ L2 norms, and their similarity post-denoising. In the last 3 columns we visualize the decomposed shared noise term $\mathcal{G}$, the L2 norm of the predicted residual term $\mathbf{h}$, and the composite noise $(\mathcal{G}+\mathbf{h})$.
@@
Figure S5: Visualization of CLIP [27] per-image denoising. We visualize all components of the per-image denoising stage. From left to right: In the first 5 columns we visualize the input image, the original noisy feature map from the model, the K-Means clusters on the original features, the L2 norm on the original features, and the similarity between the central red patch and other patches. In the next 4 columns we visualize the the denoised feature map using DVT, the denoised features’ K-means clusters, the denoised features’ L2 norms, and their similarity post-denoising. In the last 3 columns we visualize the decomposed shared noise term $\mathcal{G}$, the L2 norm of the predicted residual term $\mathbf{h}$, and the composite noise $(\mathcal{G}+\mathbf{h})$.
@@
Figure S6: Visualization of EVA02 [13] per-image denoising. We visualize all components of the per-image denoising stage. From left to right: In the first 5 columns we visualize the input image, the original noisy feature map from the model, the K-Means clusters on the original features, the L2 norm on the original features, and the similarity between the central red patch and other patches. In the next 4 columns we visualize the the denoised feature map using DVT, the denoised features’ K-means clusters, the denoised features’ L2 norms, and their similarity post-denoising. In the last 3 columns we visualize the decomposed shared noise term $\mathcal{G}$, the L2 norm of the predicted residual term $\mathbf{h}$, and the composite noise $(\mathcal{G}+\mathbf{h})$.
@@
Figure S7: Visualization of DeiT-III [36] per-image denoising. We visualize all components of the per-image denoising stage. From left to right: In the first 5 columns we visualize the input image, the original noisy feature map from the model, the K-Means clusters on the original features, the L2 norm on the original features, and the similarity between the central red patch and other patches. In the next 4 columns we visualize the the denoised feature map using DVT, the denoised features’ K-means clusters, the denoised features’ L2 norms, and their similarity post-denoising. In the last 3 columns we visualize the decomposed shared noise term $\mathcal{G}$, the L2 norm of the predicted residual term $\mathbf{h}$, and the composite noise $(\mathcal{G}+\mathbf{h})$.
@@
Figure S8: Visualization of DINOv2 with Registers [7] per-image denoising. We visualize all components of the per-image denoising stage. From left to right: In the first 5 columns we visualize the input image, the original noisy feature map from the model, the K-Means clusters on the original features, the L2 norm on the original features, and the similarity between the central red patch and other patches. In the next 4 columns we visualize the the denoised feature map using DVT, the denoised features’ K-means clusters, the denoised features’ L2 norms, and their similarity post-denoising. In the last 3 columns we visualize the decomposed shared noise term $\mathcal{G}$, the L2 norm of the predicted residual term $\mathbf{h}$, and the composite noise $(\mathcal{G}+\mathbf{h})$.
@@
Figure S9: Visualization of DINO [3] per-image denoising. We visualize all components of the per-image denoising stage. From left to right: In the first 5 columns we visualize the input image, the original noisy feature map from the model, the K-Means clusters on the original features, the L2 norm on the original features, and the similarity between the central red patch and other patches. In the next 4 columns we visualize the the denoised feature map using DVT, the denoised features’ K-means clusters, the denoised features’ L2 norms, and their similarity post-denoising. In the last 3 columns we visualize the decomposed shared noise term $\mathcal{G}$, the L2 norm of the predicted residual term $\mathbf{h}$, and the composite noise $(\mathcal{G}+\mathbf{h})$.
@@
Figure S10: Visualization of MAE [16] per-image denoising. We visualize all components of the per-image denoising stage. From left to right: In the first 5 columns we visualize the input image, the original noisy feature map from the model, the K-Means clusters on the original features, the L2 norm on the original features, and the similarity between the central red patch and other patches. In the next 4 columns we visualize the the denoised feature map using DVT, the denoised features’ K-means clusters, the denoised features’ L2 norms, and their similarity post-denoising. In the last 3 columns we visualize the decomposed shared noise term $\mathcal{G}$, the L2 norm of the predicted residual term $\mathbf{h}$, and the composite noise $(\mathcal{G}+\mathbf{h})$.
Appendix D Discussion on Limitations
Limitations.
Our approach faces some practical and theoretical challenges. On the practical front, although our method leverages parallel computing to amortize the denoising process, the time required to denoise a single image, such as one with a resolution of $518\times 518$, remains high — approximately 100 seconds. This duration may be impractical for commercial or personal users with limited access to parallel computing resources, despite the fact that we can finish denoising 10k samples within hours. Additionally, our generalizable denoisesr, trained on the last layer features of pretrained ViTs, does not remove noise in intermediate outputs. Users requiring denoised features from multiple layers might need to train distinct denoisers for different layers. From the theoretical perspective, the reasons behind the presence of these artifacts remain unclear. Integrating insights from Registers [7] with our findings could yield a more comprehensive understanding of these phenomena.
Broader Impact.